# A Survey on Text-Driven 360-Degree Panorama Generation

### v1

@@ 4.3 Quantitative Comparison
To ensure an unbiased evaluation of the generalizability of the methods, we employ an out-of-domain dataset, ODI-SR Deng et al. (2021), on which none of the models have been explicitly trained. For generating text descriptions, we utilize BLIP2 Li et al. (2023) to create textual captions for the 360-degree panoramas included in the ODI-SR dataset. These generated text prompts serve as inputs for both Text-Only Generation and Text-Driven NFoV Outpainting tasks. To simulate NFoV images, we first project the equirectangular 360-degree panoramas from the ODI-SR dataset into a cubemap format and then extract the front face of each cubemap. The original 360-degree panoramas from ODI-SR are designated as real images and used as ground truth for the computation of evaluation metrics.
The quantitative results obtained from this comparative evaluation, across the seven evaluation metrics, are presented in Tab. 4. To provide insights into the computational efficiency of each method, we also report the inference time required to generate a 1024$\times$512 360-degree panorama on a consistent machine equipped with an RTX A6000 GPU.
5 Emerging 3D Applications
Recent advances in text-driven 360-degree panorama generation Wang et al. (2024); Feng et al. (2023); Zhang et al. (2024) have catalyzed innovative methods for reconstructing 360-degree 3D scenes from textual descriptions. 360-degree panoramic images inherently capture both global contexts and geometric constraints of a scene, making them an essential intermediate representation for 3D scene generation. Consequently, recent text-driven 360-degree 3D scene generation methods Ma et al. (2024); Zhou et al. (2024a, b); Li et al. (2024); Yang et al. (2024) utilize 360-degree panorama generation to bridge the gap between text prompts and 360-degree 3D scene reconstruction.
As depicted in Fig. 2, these methods typically use a two-stage process: (1) 360-Degree Panorama Generation: generating a 360-degree panorama from the input text prompt using a fine-tuned LDM Rombach et al. (2022), and (2) 3D Scene Reconstruction: inferring a 3D representation, typically with 3D Gaussian Splatting (3DGS) Kerbl et al. (2023), from the generated panorama and corresponding multi-view perspective images. Tab. 5 provides a comparative summary.
Within this framework, emerging methods are primarily differentiated by their choice of (a) 360-degree panorama generators and their (b) strategies for extracting and utilizing 3D information. For instance, FastScene Ma et al. (2024) and HoloDreamer Zhou et al. (2024a) both employ Diffusion360 Feng et al. (2023) to generate the initial 360-degree panorama depicting a scene from a given text prompt. FastScene Ma et al. (2024) then synthesizes multi-view panoramas of this scene for specific camera poses using Coarse View Synthesis and Progressive Novel View Inpainting. With these synthesized multi-view panoramas, FastScene introduces Multi-View Projection to get their perspective views. The point clouds derived from these views are then used as input for 3DGS to reconstruct the 3D scene. HoloDreamer Zhou et al. (2024a) enhances the Diffusion360-generated panorama with two distinct ControlNet-based LDMs Zhang et al. (2023) and a super-resolution network to create a high-resolution, stylized output. Subsequently, HoloDreamer initializes 3D Gaussians using point clouds derived from a reverse equirectangular projection of the high-resolution panorama combined with its corresponding depth information. Finally, a two-stage 3DGS optimization process is developed to refine the scene rendering, resulting in the desired 3D scene reconstruction.
Furthermore, certain methods deviate from the reliance on Diffusion360 Feng et al. (2023). DreamScene360 Zhou et al. (2024b) utilizes StitchDiffusion Wang et al. (2024) to generate multiple 360-degree panorama candidates and then employs a self-refinement process to select the optimal candidate for initializing panoramic 3D Gaussians with a 3D geometric field. To facilitate visual feature correspondences between different views and maintain geometric consistencies during the 3DGS optimization process, semantic and geometric regularizations are applied. In contrast, SceneDreamer360 Li et al. (2024) uses a fine-tuned PanFusion Zhang et al. (2024) generator, coupled with a super-resolution module from Feng et al. (2023), to produce a high-resolution (6K) panorama aligned with the input text prompt. It then uses optimization-based viewpoint selection to extract multi-view images, which are subsequently used for improved point cloud initialization, ultimately leading to 3DGS-based scene reconstruction.
@@ 5 Emerging 3D Applications
Table 5: Summary of Text-Driven 360-Degree 3D Scene Generation. The column labeled ‘360-Degree PG’ indicate which text-driven 360-degree panorama generation techniques are adopted to synthesize the intermediate panoramic representation of the scene. ‘3DGS’ denotes 3D Gaussian Splatting.
Other methods explore alternative panorama generation techniques. LayerPano3D Yang et al. (2024) begins by generating four orthogonal perspective views with a fine-tuned text-to-image model Podell et al. (2024). These initial views are then combined with text-guided inpainting Rombach et al. (2022), and further processed by using a fine-tuned Diffusion360 Feng et al. (2023) model to outpaint the polar regions, resulting in a reference 360-degree panorama. To handle occlusions in complex scenes, LayerPano3D Yang et al. (2024) decomposes the reference panorama into multiple depth-based layers and utilizes a fine-tuned inpainter Zhang et al. (2024) to complete unseen content at each layer. These inpainted, layered panoramas then provide supervision for panoramic 3D Gaussian scene optimization.
6 Challenges and Future Directions
@@
Although OmniFID Christensen et al. (2024) extends FID to account for geometric fidelity in generated 360-degree panoramas, it primarily focuses on geometric aspects, leaving other crucial dimensions unaddressed. Text-to-image consistency, commonly evaluated by using the CLIP Score Radford et al. (2021), is another critical metric for assessing text-driven 360-degree panorama generation methods. However, the CLIP Score, primarily designed for conventional 2D images, relies on an image encoder predominantly trained on standard perspective images. This encoder may not fully capture the inherent distortions and spatial relationships unique to 360-degree panoramas represented by equirectangular projections. Consequently, the CLIP Score may not accurately reflect the true semantic consistency in the context of 360-degree panorama generation. Future research could explore innovative evaluation metrics that effectively assess the semantic consistency between the input prompt and the generated 360-degree panoramic image.
@@
While Diffusion360 Feng et al. (2023), which uses a super-resolution module, is among the few methods that can currently achieve a maximum resolution of 6144$\times$3072 (6K), this remains inadequate for demanding applications like VR gaming and high-fidelity 3D scene reconstruction, which often necessitate resolutions of 8K or higher to capture intricate details of landscapes and architecture. Generating such high-resolution panoramas, however, introduces significant challenges related to computational latency and memory consumption. Addressing these limitations requires the development of more efficient model architectures and optimization techniques. Promising approaches include the use of window-based operations, model pruning, quantization, and advanced neural network designs tailored to resource-intensive tasks. Moreover, the availability of high-resolution, large-scale datasets will be critical for driving progress in this direction.
@@
Existing text-driven methods, despite their ability to produce photorealistic 360-degree panoramas, often lack precise control over global semantic layout and spatial structure of the generated scene. This motivates exploring multi-modal approaches to enhance controllability. Although 360PanT Wang and Xue (2025) demonstrates panorama-to-panorama translation using auxiliary modalities like edge and segmentation maps alongside text, its outputs deviate from the standard 360°$\degree$$\times$180°$\degree$ field of view when these additional modalities are incorporated. Future research should focus on developing multi-modal techniques that effectively integrate diverse inputs (e.g. depth maps, segmentation maps, or edge maps) with text prompts to achieve fine-grained spatial control in the generated 360-degree panoramas, while ensuring strict adherence to the standard equirectangular projection format.
@@
Most text-driven 360-degree panorama generation methods are built upon latent diffusion models (LDMs) Rombach et al. (2022). While LDMs have achieved remarkable success in text-to-image synthesis, recent advancements in autoregressive models indicate promising alternative architectures. Specifically, Visual Autoregressive (VAR) models, exemplified by Infinity Han et al. (2024), have exhibited superior performance compared to the leading LDMs in standard text-to-image synthesis. This highlights an exciting avenue for future research: exploring VAR-based models for text-driven 360-degree panorama generation.
7 Conclusion
This survey has provided a comprehensive overview of the rapidly evolving field of text-driven 360-degree panorama generation. We began by introducing two primary representation methods of 360-degree panoramas, along with widely used datasets, and key evaluation metrics in this domain. Subsequently, we presented an in-depth discussion of prevalent methods for text-driven 360-degree panorama generation and their applications in generating 360-degree 3D scene. Despite the significant progress achieved in this field, several challenges remain. To address these challenges, we have articulated promising directions for future research.

### v2

@@ V Emerging 3D Applications
TABLE V: Summary of Text-Driven 360-Degree 3D Scene Generation. The column labeled ‘360-Degree PG’ indicate which text-driven 360-degree panorama generation techniques are adopted to synthesize the intermediate panoramic representation of the scene. ‘3DGS’ denotes 3D Gaussian Splatting.
@@
LayerPano3D SIGGRAPH 2025 Diffusion360 & PanFusion 3DGS
As depicted in Fig. 8, these methods typically use a two-stage process: (1) 360-Degree Panorama Generation: generating a 360-degree panorama from the input text prompt using a fine-tuned LDM [63], and (2) 3D Scene Reconstruction: inferring a 3D representation, typically with 3D Gaussian Splatting (3DGS) [26], from the generated panorama and corresponding multi-view perspective images. Tab. V provides a comparative summary.
Within this framework, emerging methods are primarily differentiated by their choice of (a) 360-degree panorama generators and their (b) strategies for extracting and utilizing 3D information. For instance, FastScene [31] and HoloDreamer [30] both employ Diffusion360 [58] to generate the initial 360-degree panorama depicting a scene from a given text prompt. FastScene [31] then synthesizes multi-view panoramas of this scene for specific camera poses using Coarse View Synthesis and Progressive Novel View Inpainting. With these synthesized multi-view panoramas, FastScene introduces Multi-View Projection to get their perspective views. The point clouds derived from these views are then used as input for 3DGS to reconstruct the 3D scene. HoloDreamer [30] enhances the Diffusion360-generated panorama with two distinct ControlNet-based LDMs [37] and a super-resolution network to create a high-resolution, stylized output. Subsequently, HoloDreamer initializes 3D Gaussians using point clouds derived from a reverse equirectangular projection of the high-resolution panorama combined with its corresponding depth information. Finally, a two-stage 3DGS optimization process is developed to refine the scene rendering, resulting in the desired 3D scene reconstruction.
Furthermore, certain methods deviate from the reliance on Diffusion360 [58]. DreamScene360 [33] utilizes StitchDiffusion [57] to generate multiple 360-degree panorama candidates and then employs a self-refinement process to select the optimal candidate for initializing panoramic 3D Gaussians with a 3D geometric field. To facilitate visual feature correspondences between different views and maintain geometric consistencies during the 3DGS optimization process, semantic and geometric regularizations are applied. In contrast, SceneDreamer360 [32] uses a fine-tuned PanFusion [59] generator, coupled with a super-resolution module from [58], to produce a high-resolution (6K) panorama aligned with the input text prompt. It then uses optimization-based viewpoint selection to extract multi-view images, which are subsequently used for improved point cloud initialization, ultimately leading to 3DGS-based scene reconstruction.
Other methods explore alternative panorama generation techniques. LayerPano3D [27] begins by generating four orthogonal perspective views with a fine-tuned text-to-image model [21]. These initial views are then combined with text-guided inpainting [63], and further processed by using a fine-tuned Diffusion360 [58] model to outpaint the polar regions, resulting in a reference 360-degree panorama. To handle occlusions in complex scenes, LayerPano3D [27] decomposes the reference panorama into multiple depth-based layers and utilizes a fine-tuned inpainter [59] to complete unseen content at each layer. These inpainted, layered panoramas then provide supervision for panoramic 3D Gaussian scene optimization.
VI Challenges and Future Directions
@@
Although OmniFID [51] extends FID to account for geometric fidelity in generated 360-degree panoramas, it primarily focuses on geometric aspects, leaving other crucial dimensions unaddressed. Text-to-image consistency, commonly evaluated by using the CLIP Score [52], is another critical metric for assessing text-driven 360-degree panorama generation methods. However, the CLIP Score, primarily designed for conventional 2D images, relies on an image encoder predominantly trained on standard perspective images. This encoder may not fully capture the inherent distortions and spatial relationships unique to 360-degree panoramas represented by equirectangular projections. Consequently, the CLIP Score may not accurately reflect the true semantic consistency in the context of 360-degree panorama generation. Future research could explore innovative evaluation metrics that effectively assess the semantic consistency between the input prompt and the generated 360-degree panoramic image.
@@
While Diffusion360 [58], which uses a super-resolution module, is among the few methods that can currently achieve a maximum resolution of 6144$\times$3072 (6K), this remains inadequate for demanding applications like VR gaming and high-fidelity 3D scene reconstruction, which often necessitate resolutions of 8K or higher to capture intricate details of landscapes and architecture. Generating such high-resolution panoramas, however, introduces significant challenges related to computational latency and memory consumption. Addressing these limitations requires the development of more efficient model architectures and optimization techniques. Promising approaches include the use of window-based operations, model pruning, quantization, and advanced neural network designs tailored to resource-intensive tasks. Moreover, the availability of high-resolution, large-scale datasets will be critical for driving progress in this direction.
@@
Existing text-driven methods, despite their ability to produce photorealistic 360-degree panoramas, often lack precise control over global semantic layout and spatial structure of the generated scene. This motivates exploring multi-modal approaches to enhance controllability. Although 360PanT [28] demonstrates panorama-to-panorama translation using auxiliary modalities like edge and segmentation maps alongside text, its outputs deviate from the standard 360°$\degree$$\times$180°$\degree$ field of view when these additional modalities are incorporated. Future research should focus on developing multi-modal techniques that effectively integrate diverse inputs (e.g. depth maps, segmentation maps, or edge maps) with text prompts to achieve fine-grained spatial control in the generated 360-degree panoramas, while ensuring strict adherence to the standard equirectangular projection format.
@@
Most text-driven 360-degree panorama generation methods are built upon latent diffusion models (LDMs) [63]. While LDMs have achieved remarkable success in text-to-image synthesis, recent advancements in autoregressive models indicate promising alternative architectures. Specifically, Visual Autoregressive (VAR) models, exemplified by Infinity [2], have exhibited superior performance compared to the leading LDMs in standard text-to-image synthesis. This highlights an exciting avenue for future research: exploring VAR-based models for text-driven 360-degree panorama generation.
VII Conclusion
This survey has provided a comprehensive overview of the rapidly evolving field of text-driven 360-degree panorama generation. We began by introducing primary representation methods of 360-degree panoramic images, along with widely used datasets, and key evaluation metrics in this domain. Subsequently, we presented an in-depth discussion of prevalent methods for text-driven 360-degree panorama generation and their applications in generating 360-degree 3D scene. Despite the significant progress achieved in this field, several challenges remain. To address these challenges, we have articulated promising directions for future research.